# Q-Net: Transferable Queue Length Estimation via Kalman-based Neural Networks

### v1

@@ 5.1 Learned Kalman gain
  • The measurement difference $\Delta\tilde{\mathbf{y}}_{t}=\mathbf{y}_{t}-\mathbf{y}_{t-1}$;
  • The innovation difference $\Delta{\mathbf{y}}_{t}=\mathbf{y}_{t}-\mathbf{\hat{y}}_{t|t-1}$;
  • The forward evolution difference $\Delta\tilde{x}_{t}=\hat{x}_{t|t}-\hat{x}_{t-1|t-1}$;
  • The forward update difference $\Delta\hat{x}_{t}=\hat{x}_{t|t}-\hat{x}_{t|t-1}$.
The forward evolution difference $\Delta\tilde{x}_{t}$ represents the change between two consecutive a posteriori state estimates, while the forward update difference $\Delta\hat{x}_{t}$ captures the correction from the a priori to the a posteriori estimate at the same time step. However, since the a posteriori estimate $\hat{x}_{t|t}$ is not available before the Kalman gain is computed, we use the previous values $\Delta\tilde{x}_{t-1}$ and $\Delta\hat{x}_{t-1}$ during inference. In this setup, measurement difference $\Delta\tilde{\mathbf{y}}_{t}$ and forward evolution difference $\Delta\tilde{x}_{t-1}$ capture the underlying state evolution dynamics, whereas the innovation difference $\Delta{\mathbf{y}}_{t}$ and the forward update difference $\Delta\hat{x}_{t-1}$ reflect the uncertainty associated with the current state estimate.
Figure 4: KalmanNet architecture for dynamic Kalman gain estimation. The top panel illustrates the overall KalmanNet pipeline, which takes in four temporal input features (measurement difference $\Delta\tilde{\mathbf{y}}_{t}$, innovation difference $\Delta{\mathbf{y}}_{t}$, forward evolution difference $\Delta\tilde{x}_{t-1}$, forward update difference $\Delta\hat{x}_{t-1}$) to estimate the Kalman gain. Intermediate representations ($\hat{Q}_{t},\hat{\Sigma}_{t},\hat{\mathbf{S}}_{t}$ and $\hat{\Sigma}_{t}^{h}$) are computed through sequential submodules. The lower panels detail the submodules a, b, c, d, and e.
As shown in Figure 4, KalmanNet can be globally divided into five submodules:
@@
    Process noise covariance module: This module estimates the process noise covariance matrix $\hat{Q}_{t}$ based on the forward update difference $\Delta\hat{x}_{t-1}$. The design is motivated by Equations (1) and (3), which suggest that evolution noise influences the predicted state estimate. The estimation follows:
@@
    The fully connected (FC) layers transform the input into a higher-dimensional feature representation. The gated recurrent unit (GRU) maintains a hidden state that is updated at each time step through its gating mechanism, enabling $\hat{Q}_{t}$ to be estimated based on the complete history of forward update differences up to the current time rather than using fixed noise models.
@@
    State covariance module: This module estimates the second-order moments of the model state, denoted as $\hat{\Sigma}_{t}$. According to Equation (4), the a posteriori state covariance depends on the process noise covariance $\hat{Q}_{t}$. Additionally, the forward evolution difference $\Delta\tilde{x}_{t-1}$ is included to capture the queue length evolution over time and provides system dynamics information that is not captured by process noise alone. The estimation is performed as follows:
@@
    where $\hat{\Sigma}_{t}^{h}$ denotes the hidden state of the GRU, later refined in the hidden state update module, and Concat denotes concatenation. The GRU’s gating mechanism selectively retains relevant historical queue length covariance patterns while discarding outdated information.
@@
    Measurement covariance module: This module estimates the measurement covariance matrix $\hat{\mathbf{S}}_{t}$. From Equation (6), $\hat{\mathbf{S}}_{t}$ is related to the innovation difference $\Delta\mathbf{y}_{t}$, which reflects speed prediction errors, and the estimated state covariance $\hat{\Sigma}_{t}$, which indicates queue length uncertainty. We also include the measurement difference $\Delta\tilde{\mathbf{y}}_{t}$ as input, as it captures temporal variations in raw aFCD speed measurement:
@@
    The GRU uses a reset gate to determine how much of the past measurement covariance pattern remains valid, and an update gate to control how much the current uncertainty from speed prediction errors, queue length uncertainty, and speed variations should influence the measurement covariance estimate.
@@
    Kalman gain module: This module computes the Kalman gain $\mathcal{K}_{t}$ based on the queue length and speed measurement covariances, consistent with the input of Kalman gain calculation in traditional Kalman filters:
@@
    Equations (3)-(8) show the recursive nature of Kalman filtering, where each Kalman gain updates the state estimate, which subsequently influences the next gain computation. This recursive nature of Kalman gain calculation is reflected in the design of the state covariance module (for $\hat{\Sigma}_{t}$) and measurement covariance module (for $\hat{\mathbf{S}}_{t}$), both of which use GRUs to capture temporal dependencies in the filtering process.
@@
In this way, KalmanNet dynamically tracks the underlying statistical moments associated with state evolution and measurements. The Kalman gain is computed adaptively, leveraging both memory of past information and current predicted state and measurement data. For additional implementation details, please refer to Revach et al. (2022).
@@ 5.2 Adaptation for spatial transferability
Recall that a segment is the spatial unit over which aFCD aggregates speed measurements. While KalmanNet offers flexibility in nonlinear Kalman filtering without prior noise assumptions, it requires fixed-dimensional measurement vectors corresponding to the number of segments. Since obtaining ground truth data for training can be difficult or costly, transferring models across sections with varying segment counts is essential, necessitating spatial transferability adaptation.
To address this, we impose that the measurement input to KalmanNet remains fixed-dimensional regardless of the number of road segments, enabling Q-Net to be trained on one section and applied to others with spatial transferability. Specifically, we assume that the covariance entries for each segment depend only on the segment itself and its immediate neighbors (previous and next segments), which is justified by the short 10-second filtering interval (ground truth frequency). As shown in Figure 5, we adopt a local measurement grouping strategy, where each measurement group for segment $1<i<N$ is defined as:
@@
We exclude boundary segments ($i=1$ and $i=N$) from the update process, as queue dynamics are primarily reflected by intermediate segments. The first segment near the stop line can be erroneous due to localized queue dissipation patterns, while the last segment provides limited information since intermediate segments repeatedly confirm the presence of a long queue when it extends to the section boundary.
These grouped features are input to KalmanNet, which produces a group-specific Kalman gain $\mathcal{K}^{i}_{t}$. Note that each group is not expected to share the same covariance matrix; instead, the model learns a parameterized function that is shared across groups, enabling adaptive estimation of the covariance structure for each local neighborhood. The state estimate is then updated by aggregating the additive contributions of all groups, preserving the structure of the classical Kalman filter:
$\hat{x}_{t|t}=\hat{x}_{t|t-1}+\sum_{i=2}^{N-1}\mathcal{K}^{i}_{t}\Delta\bar{\mathbf{y}}^{i}_{t}$ (23)
Although individual measurements may appear in multiple overlapping groups, KalmanNet learns appropriate group-specific gains that implicitly balance their influence. This allows the network to determine how much each local group should affect the global state estimate. Moreover, this grouping strategy enables KalmanNet to generalize across road sections with varying segmentations. By focusing on the current segment and its adjacent segments, the model captures local neighborhood covariances, enhancing both robustness and transferability.
@@ 6 Numerical results
In this section, we validate our method Q-Net using a real-world case study from Rotterdam, the Netherlands. The dataset spans 18 days (from 6:00 AM to 8:00PM), from 2023-11-10 to 2023-11-27, comprising 12 weekdays and 6 weekend days. We randomly split the data into training, validation, and test sets while ensuring a balanced representation of weekdays and weekends. Specifically, 2023-11-11 (Saturday) and 2023-11-20 (Monday) are used for validation, while 2023-11-14 (Tuesday), 2023-11-18 (Saturday), and 2023-11-22 (Wednesday) are designated as test days.
We present the experimental setup in Subsection 6.1, evaluate model performance against baselines in Subsection 6.2, assess transferability in Subsection 6.3, demonstrate real-time applicability in Subsection 6.4, and analyze component contributions through ablation studies in Subsection 6.5.
6.1 Experiment setup
Q-Net is trained end-to-end in a supervised fashion and optimized using the Adam optimizer. When aFCD data is missing, the missing value is imputed by the previous value of the same segment. The loss function is defined as the Root Mean Square Error (RMSE) between the estimated queue length $\hat{x}_{t|t}$ and the ground truth queue length $x_{t}$ measured by radar data. Given a sequence of $T$ time steps, the loss is computed as:
$\mathcal{L}=\sqrt{\frac{\sum_{t=0}^{T-1}(x_{t}-\hat{x}_{t|t}^{2})^{2}}{T}}$ (24)
During training, the clipping functions in Equations (15) and (17) are disabled to ensure uninterrupted gradient backpropagation. These functions are applied only during validation and testing to enforce realistic queue length constraints.
In this study, each time series spans from 6:00 AM to 8:00 PM. With a time resolution of 10 seconds, this results in 5,040 time steps per sequence. Due to the long sequence length, accumulated prediction errors during the early stages of training may hinder effective learning. To mitigate this, we divide the sequence into shorter windows of 10 minutes for neural network training and model weight updates. The system state at the beginning of each window is initialized to zero if it is the first window of the day; otherwise, it is initialized using the a posteriori state estimate from the preceding window.
Q-Net is computationally efficient, comprising only 952 trainable parameters and demonstrating fast training performance. When trained on an Apple M2 Pro chip using 13 days of data, each epoch takes approximately 40 seconds. The cutoff frequencies of the Fourier-based bandpass filter can be empirically determined by visual inspection of the filtered data. The KalmanNet implementation follows that of Revach et al. (2022) [^1] https://github.com/KalmanNet/KalmanNet_TSP, with the learning rate set to 0.001.
@@ 6.1 Experiment setup
We compare our model against baseline approaches. First, we review prior work using similar datasets, namely loop detector counts and/or aFCD speed data. For example, Luo et al. (2022) combines aFCD with loop detector data but also requires traffic light event data. Models relying solely on loop detector data are typically designed for cycle-wise maximum queue length estimation (Liu et al., 2009; Abewickrema et al., 2023; Sengupta et al., 2021; An et al., 2018), or assume detectors placed at the entrance of the section with fully observed and conserved traffic flow (Heshami and Kattan, 2021; Lee et al., 2019; Yang et al., 2019; Karnati et al., 2021). These assumptions are not compatible with our study context, which includes a large unobserved traffic flow and limited loops. As a result, we select two simple yet representative methods that rely only on aFCD as our baselines:
  • OSD (Altıntaşı, 2018), which detects the end of the queue based on observed speed drops where the upstream speed is higher than 16 km/h and downstream speed is lower than 16 km/h.
  • ISC (Dinh et al., 2014), which constructs synthetic trajectories from aFCD and estimates the queue using iso-speed contours. Speed thresholds are selected to be 14 km/h, 16 km/h, 18 km/h, 20 km/h, and 22 km/h.

### v2

@@ 5.1 Learned Kalman gain
Figure 4: KalmanNet architecture for dynamic Kalman gain estimation. The top panel illustrates the overall KalmanNet pipeline, which takes in four temporal input features (forward evolution difference $\Delta\tilde{x}_{t-1}$, forward update difference $\Delta\hat{x}_{t-1}$, measurement difference $\Delta\tilde{\mathbf{y}}_{t}$, innovation difference $\Delta{\mathbf{y}}_{t}$) to estimate the Kalman gain. Intermediate representations ($\hat{Q}_{t},\hat{\Sigma}_{t},\hat{\mathbf{S}}_{t}$ and $\hat{\Sigma}_{t}^{h}$) are computed through sequential submodules. The lower panels detail the submodules a, b, c, d, and e.
@@
The $\textit{forward evolution difference: }\quad\Delta\tilde{x}_{t}=\hat{x}_{t|t}-\hat{x}_{t-1|t-1};$ (22)
The $\textit{forward update difference: }\quad\Delta\hat{x}_{t}=\hat{x}_{t|t}-\hat{x}_{t|t-1};$ (23)
The $\textit{measurement difference: }\quad\Delta\tilde{\mathbf{y}}_{t}=\mathbf{y}_{t}-\mathbf{y}_{t-1};$ (24)
The $\textit{innovation difference: }\quad\Delta{\mathbf{y}}_{t}=\mathbf{y}_{t}-\mathbf{\hat{y}}_{t|t-1}.$ (25)
The forward evolution difference $\Delta\tilde{x}_{t}$ represents the change between two consecutive a posteriori state estimates, while the forward update difference $\Delta\hat{x}_{t}$ captures the correction from the a priori to the a posteriori estimate at the same time step. However, since the a posteriori estimate $\hat{x}_{t|t}$ is not available before the Kalman gain is computed, we use the previous values $\Delta\tilde{x}_{t-1}$ and $\Delta\hat{x}_{t-1}$ during inference. In this setup, forward evolution difference $\Delta\tilde{x}_{t-1}$ and measurement difference $\Delta\tilde{\mathbf{y}}_{t}$ capture the underlying state evolution dynamics, whereas the forward update difference $\Delta\hat{x}_{t-1}$ and the innovation difference $\Delta{\mathbf{y}}_{t}$ reflect the uncertainty associated with the current state estimate.
Recall that $\Sigma_{t}$ and $\mathbf{S}_{t}$ denote the covariance matrices of the state $x_{t}$ and the measurements $\mathbf{y}_{t}$, while $Q_{t}$ and $\mathbf{R}_{t}$ represent the process and measurement noise covariances. Let $\hat{(\cdot)}$ denote the estimated quantities and ${(\cdot)}^{h}$ indicate latent representations. As shown in Figure 4, KalmanNet can be globally divided into five submodules:
The process noise covariance module estimates the process noise covariance $\hat{Q}_{t}$ based on the forward update difference $\Delta\hat{x}_{t-1}$. Both $\hat{Q}_{t}$ and the forward evolution difference $\Delta\tilde{x}_{t-1}$ are then fed into the state covariance module to produce $\hat{\Sigma}_{t}$. The latent representation $\hat{\Sigma}_{t}^{h}$ is later refined by the hidden state covariance module. Next, $\hat{\Sigma}_{t}$, the innovation difference $\Delta\mathbf{y}_{t}$, and the measurement difference $\Delta\tilde{\mathbf{y}}_{t}$ are provided to the measurement covariance module to compute $\hat{\mathbf{S}}_{t}$. Finally, $\hat{\Sigma}_{t}$ and $\hat{\mathbf{S}}_{t}$ are passed to the Kalman gain module, which outputs the Kalman gain $\mathcal{K}_{t}$. All covariance matrices are represented in vectorized form throughout the network, which simplifies operations such as concatenation and allows the network to learn each entry without explicit matrix structure constraints.
@@
    Process noise covariance module: The design is motivated by Equations (1) and (3), which suggest that evolution noise influences the predicted state estimate. The estimation follows:
@@
    The fully connected (FC) layers map the input into a higher-dimensional feature space. The gated recurrent unit (GRU) maintains a hidden state that is updated at each time step through its gating mechanism, allowing $\hat{Q}_{t}$ to reflect the complete history of forward update differences rather than a static noise model.
@@
    State covariance module: This module estimates $\hat{\Sigma}_{t}$ using the process noise covariance $\hat{Q}_{t}$ (per Equation (4)) and the forward evolution difference $\Delta\tilde{x}_{t-1}$, which captures queue length evolution and system dynamics not reflected in process noise alone:
@@
    where $\hat{\Sigma}_{t}^{h}$ denotes the GRU hidden state (later refined in module e) and Concat denotes concatenation. The GRU selectively retains relevant historical covariance patterns while discarding outdated information.
@@
    Measurement covariance module: Throughout KalmaNet, we target the measurement covariance $\mathbf{S}_{t|t-1}$ and denote its learned surrogate as $\hat{\mathbf{S}}_{t}$ for brevity. From Equation (6), $\hat{\mathbf{S}}_{t}$ depends on the innovation difference $\Delta\mathbf{y}_{t}$ (speed prediction errors) and $\hat{\Sigma}_{t}$ (queue length uncertainty). We additionally include the measurement difference $\Delta\tilde{\mathbf{y}}_{t}$ to capture temporal variations in raw aFCD speed measurements, and the GRU gates control how past measurement covariance patterns and current uncertainties influence the estimate.:
@@
    Kalman gain module: This module computes the Kalman gain $\mathcal{K}_{t}$ based on the queue length and speed measurement covariances, consistent with the inputs used by Kalman gain computation (state and measurement uncertainties):
@@
    The recursive nature of Kalman gain calculation (Equations (3)-(10)), where each Kalman gain updates the state estimate and subsequently influences the next gain computation, is reflected through the GRU-based temporal dependencies in modules b and c.
@@
In this way, KalmanNet dynamically tracks the underlying statistical moments associated with state evolution and measurements. The Kalman gain is computed adaptively, leveraging both memory of past information and current predicted state and measurement data:
$\mathcal{K}_{t}=\text{KalmanNet}(\Delta\tilde{x}_{t},\Delta\hat{x}_{t},\Delta\tilde{\mathbf{y}}_{t},\Delta{\mathbf{y}}_{t}).$ (31)
For additional implementation details, please refer to Revach et al. (2022).
@@ 5.2 Adaptation for spatial transferability
A segment is the spatial unit over which aFCD aggregates speed measurements. While KalmanNet offers flexibility in nonlinear Kalman filtering without prior noise assumptions, it requires fixed-dimensional measurement vectors corresponding to the number of segments. Since obtaining ground truth data for training can be difficult or costly, transferring models across sections with varying segment counts is essential, necessitating spatial transferability adaptation.
To address this, we ensure that the measurement input to KalmanNet has fixed dimensionality regardless of the number of road segments, enabling Q-Net to be trained on one section and transferred to others. Specifically, we assume that covariance entries for each segment depend only on the segment itself and its immediate neighbors (previous and next segments), which is justified by the short 10-second filtering interval (ground truth frequency). As shown in Figure 5, we adopt a local measurement grouping strategy where each measurement group for segment $1<i<N$ comprises:
@@
We exclude boundary segments ($i=1$ and $i=N$) because queue dynamics are primarily reflected by intermediate segments. The first segment near the stop line can be erroneous due to localized queue dissipation patterns, while the last segment provides limited information since intermediate segments repeatedly confirm the presence of a long queue when it extends to the section boundary.
These grouped features are fed to the (c) measurement covariance module of KalmanNet. This is where our approach departs from a classical Kalman filter. In a scalar-state setup, the classical update (Equation (8)) is effectively a single weighted sum, where each innovation $\Delta y_{t}^{j}$ is multiplied by a gain component $(\mathcal{K}_{t})_{1,j}$ computed from global covariance matrices. Q-Net, however, does not learn one global $\mathcal{K}_{t}$. It learns multiple, local gain vectors $\mathcal{K}^{i}_{t}$ for each group $i$:
$\mathcal{K}_{t}^{i}$ $=\text{KalmanNet}(\Delta\tilde{x}_{t},\Delta\hat{x}_{t},\Delta\breve{\mathbf{y}}^{i}_{t},\Delta{\mathbf{\bar{y}}}^{i}_{t}),$ $\text{where }\mathcal{K}^{i}_{t}\in\mathbb{R}^{1\times 3},$ (32)
$\hat{x}_{t|t}$ $=\hat{x}_{t|t-1}+\sum_{i=2}^{N-1}\mathcal{K}^{i}_{t}\Delta\bar{\mathbf{y}}^{i}_{t},$ $\text{where }\Delta{\mathbf{\bar{y}}}^{i}_{t}\in\mathbb{R}^{3}.$ (33)
Each group shares the same parameterized function. Since a single measurement innovation $\Delta y_{t}^{j}$ appears in three groups (j1j-1, $j$, and $j+1$), its total effective gain is implicitly the sum of components from three adjacent group gains ($\mathcal{K}^{j-1}_{t}$, $\mathcal{K}^{j}_{t}$, and $\mathcal{K}^{j+1}_{t}$). While this is not a globally optimal update in the classical sense, this design localizes the gain calculation. This localization makes the architecture independent of the total section segment number ($N$) and thereby achieving spatial transferability.
5.3 Training
Q-Net is parameterized to learn the Kalman gain and is trained in an end-to-end manner. The loss function is the Root Mean Square Error (RMSE) between the estimated queue length $\hat{x}_{t|t}$ and the ground truth queue length $x_{t}$ measured by radar data. Given a sequence of $T$ time steps, the loss is computed as:
$\mathcal{L}=\sqrt{\frac{1}{T}\sum_{t=0}^{T-1}(x_{t}-\hat{x}_{t|t})^{2}}$ (34)
During training, the projections in Equations (19) and (21) are disabled to ensure uninterrupted gradient flow. These constraints are applied only during validation and testing to enforce physical queue length bounds.
@@ 6 Numerical results
In this section, we evaluate Q-Net, focusing on its empirical performance (Section 6.1), spatial transferability (Section 6.2), real-time capabilities (Section 6.3), and the role of its internal components (Section 6.4).
Case study. We consider a real-world case study from Rotterdam, the second-largest city in the Netherlands, characterized by a dense urban road network and significant traffic congestion. The dataset spans 16 days (from 6:00 AM to 8:00 PM), comprising 12 weekdays and 4 weekend days. As a preprocessing step, missing aFCD data were imputed using the most recent available value within the same segment.
To evaluate our method, we first filter out road segments lacking sufficient data availability, either due to contractual restrictions with the municipality or significant missing values. This ensures the analysis focuses on queue estimation rather than data imputation. Among the remaining candidates, we select N1-IN (Figure 6). This major urban road section is of practical importance as it directly connects to the Maastunnel, a critical focus of the municipality’s traffic management. This section presents a challenging scenario due to considerable unobserved traffic flow exiting toward a ring interchange.
@@
Baselines. We compare Q-Net against relevant baseline approaches. Many prior works are not directly comparable due to incompatible assumptions. For example, models relying solely on loop detector data often estimate cycle-wise maximum queue length (Liu et al., 2009; Abewickrema et al., 2023; Sengupta et al., 2021a; An et al., 2018) or assume fully observed and conserved traffic flow (Heshami and Kattan, 2021; Lee et al., 2019; Yang et al., 2019; Karnati et al., 2021). These assumptions are not compatible with our study context, which features large unobserved traffic flow and limited loops. Furthermore, cycle-wise methods estimate per-lane, per-signal-group maximum queues, whereas our ground truth is per-section every 10 seconds, making direct evaluation across granularities difficult.
Given these constraints, we selected three baselines. The first two are simple yet representative methods using only aFCD: OSD (Altıntaşı, 2018), which detects the end of the queue based on observed speed drops (upstream >> 16 km/h, downstream << 16 km/h); and ISC (Dinh et al., 2014), which constructs synthetic trajectories from aFCD and estimates the queue using iso-speed contours (with thresholds from 14 to 22 km/h). The third, Tree-RPROP (Luo et al., 2022), serves as a strong, data-fusing baseline. It combines aFCD, loop detector, and traffic light event data, employing a decision tree to detect queue presence and a resilient propagation neural network to estimate its length. For this baseline, we adopted the same setup as the original paper.
Figure 7: Six-fold data split: for each split, samples from both weekdays and weekends were randomly selected to form the training, validation, and test sets.
Q-Net implementation and hyperparameters. The cutoff frequencies of the Fourier-based band-pass filter determine the range of temporal variations preserved in the signal. After fine-tuning on the validation set, the lower and upper cutoff frequencies were set to 1/360 and 1/30 cycles per step. Given that the system time unit is 10s, this corresponds to a pass-band for signal periods between 300s (30 steps) and 3600s (360 steps).
The KalmanNet implementation follows that of Revach et al. (2022) [^2] https://github.com/KalmanNet/KalmanNet_TSP. The Adam optimizer is used for training with learning rate 0.002. Q-Net comprises 952 trainable parameters, and each training epoch requires approximately 40 seconds on an Apple M2 Pro chip.
Training strategy. At a 10s resolution, each daily series comprises 5,040 time steps. Such long sequences can cause accumulated prediction errors to hinder effective learning during the early stages of training. To mitigate this, we divide the sequence into shorter 10-minute windows for training. The state is initialized to zero for the first window of each day. For all subsequent windows, the initial state is set to the a posteriori state estimate from the end of the preceding window, which mimics practical application where the true state is unobservable.
Evaluation protocol. Due to the limited data, we perform six independent train–validation–test splits to ensure a robust evaluation. The model was trained and tested independently on each split. For each split, we randomly sample 11 days for training, 2 for validation, and 3 for testing, ensuring a mix of weekday and weekend samples (Figure 7). This process yields a total of 18 test-day evaluations and covers every day in the complete dataset for testing at least once. Details of the day-wise split are provided in the Appendix A.
Models incorporating randomness are trained using five random seeds. Performance is evaluated using Root Mean Square Error (RMSE) in meters, Mean Absolute Error (MAE) in meters, and Mean Absolute Percentage Error (MAPE) in percent (%). The metrics are computed by concatenating predictions and ground truths across all test splits. To prevent skew from large relative errors at near-zero values, MAPE is computed only for instances where the ground-truth queue length exceeds 10 meters (71.73% of the test set). Recognizing that traffic dynamics vary across time periods, metrics are reported for the entire day, with additional analysis provided for peak periods when relevant. The morning (07:00–09:00) and afternoon (16:00–18:00) peak hours are defined following Dutch Ministry of Infrastructure guidelines (Mobiliteitsscan, ).